Circular No. 8863

Central Bureau for Astronomical Telegrams INTERNATIONAL ASTRONOMICAL UNION

Mailstop 18, Smithsonian Astrophysical Observatory, Cambridge, MA 02138, U.S.A.

IAUSUBS@CFA.HARVARD.EDU or FAX 617-495-7231 (subscriptions) CBAT@CFA.HARVARD.EDU (science)

        URL http://www.cfa.harvard.edu/iau/cbat.html ISSN 0081-0304

Phone 617-495-7440/7244/7444 (for emergency use only)

COMET C/2007 Q1 (GARRADD) G. J. Garradd reports his discovery of a slightly diffuse comet on CCD

images taken in poor seeing with the 0.5-m Uppsala Schmidt telescope in the course of the Siding Spring Survey (discovery observation tabulated below). Following posting on the Minor Planet Center’s ‘NEOCP’ webpage, several other CCD astrometrists have commented on the object’s cometary appearance. N. Teamo and S. Hönig (0.41-m f/8 Ritchey-Chrétien telescope at Punaauia, Tahiti) found a 6′′ diffuse coma and a fan-shapped 20′′-long tail in p.a. ≈ 200^{o}, spanning ≈ 50^{o} of p.a., on Aug. 22.5. L. Ligustri (Talmassons, Udine, Italy, 0.35-m f/5 reflector) writes that fifteen 180-s co-added exposures from Aug. 24.0 show a coma diameter of ∼ 12′′ with an elongation in p.a. 215^{o}. L. Buzzi and F. Luppi (Varese, Italy, 0.60-m f/4.6 reflector) note that their images on Aug. 24.05–24.10 show a condensed 10′′ coma with a possible extension 6′′–8′′ long in p.a. 225^{o}. A. C. Gilmore (Mt. John 1.0-m f/7.7 reflector) adds that his images from Aug. 24.5 show a small, slightly condensed spot slightly larger than nearby star images of similar brightness.

2007 UT  α_{2000}  δ_{2000}  Mag.
Aug. 21.65039  1^{h}33^{m}57.^{s}81  −17^{o}49′48′′.2  19.4

The available astrometry, the following preliminary parabolic orbital elements, and an ephemeris appear on MPEC 2007-Q20.

T = 2006 Dec. 6.785 TT  ω = 276.^{o}821
Ω = 6.578 }2000.0
q = 2.77670 AU  i = 79.553

V458 VULPECULAE = NOVA VULPECULAE 2007 N. N. Samus, Institute of Astronomy, Russian Academy of Sciences,

informs us that this nova (cf. IAUC 8861, 8862; CBETs 1029, 1038) has been given the designation V458 Vul.

Visual magnitude estimates by C. Labordena, Castelló, Spain: Aug. 12.856 UT, 9.1; 19.853, 8.6; 21.890, 10.1. CCD magnitudes: Aug. 22.483, R = 9.51 ± 0.05 (E. Forné, L’Ampolla, Spain); 22.491, V = 10.48 ± 0.07 (Forné); 23.003, B = 11.72 ± 0.06 (Forné); 23.828, R = 9.10 ± 0.16 (A. Ardanuy, Sabadell, Spain); 23.835, V = 11.02 ± 0.09 (Ardanuy); 23.839, B = 11.45 ± 0.11 (Ardanuy).

2007 August 24  \begin{matrix}c \\ ◯ \end{matrix} Copyright 2007 CBAT  Daniel W. E. Green